# Deep learning joint extremes of metocean variables using the SPAR model

### v1

@@ 3.1 Angular modelling
Estimation of densities on the hypersphere $\mathbb{S}^{d-1}$ is part of a discipline known as directional statistics [mardia2000, ley2017modern, Pewsey2021]. The key difference from estimation of densities on $\mathbb{R}^{d}$ is that the surface of the hypersphere is periodic and bounded, and so distributions defined on $\mathbb{S}^{d-1}$ must conserve these constraints. Various parametric and non-parametric approaches have been developed for estimating densities on the hypersphere, which are directly analogous to approaches used in Euclidean space. These include kernel density estimation [hall1987kernel, bai1989kernel], mixture models [mooney2003fitting, fu2008modified], and spline-based methods [ferreira2008directional]. See [Pewsey2021] for a recent review of non-parametric approaches.
Previous uses of SPAR in two dimensions adopted a kernel density method for estimating the angular density. Although this approach can be applied in higher dimensions, simulation from the estimated density can be very slow, as it requires transformation to hyperspherical coordinates and numerical integration to obtain conditional distributions. The use of parametric mixture models instead results in faster simulation. However, for the application described in Section 4, we found that a mixture of $d$-variate von Mises distributions [hornik2014movmf] was not flexible enough to capture the complex structure in the angular data, even with hundreds of mixture components. Moreover, spline-based approaches have only been proposed for $d\leq 3$, and are therefore not applicable for the setting we consider.
As noted in Section 2, an estimate of the angular density function $f_{\mathbf{W}}$ is required to approximate the joint density of the random vector $\mathbf{X}$. However, to simulate from the SPAR model, one only requires a means of simulating from the angular variable $\mathbf{W}$; an explicit form for the density is not a necessity. Consequently, for our approach, we opt to sample the angular component empirically, i.e., we draw from the observed sample $\{\mathbf{w}_{1},\ldots,\mathbf{w}_{n}\}$ with replacement. Since the angular variable lies in a bounded domain (the surface of the hypersphere), values of $\mathbf{W}$ are never ‘extreme’. Therefore, given a sufficiently large sample size $n$, empirical resampling should provide a representative sample of the angular variable $\mathbf{W}$ for all regions of $\mathbb{S}^{d-1}$ where this variable has probability mass. Providing that we can also simulate from the conditional radial variable, $R|\mathbf{W}=\mathbf{w},$ we can generate new, representative data in the joint tail of $\mathbf{X}$. However, as discussed further in Section 4, in regions where angular observations are sparse, resampling does not give us a method of simulating unobserved, but plausible angles. Alternative methods of modelling angular density will be examined in future work.
@@ 3.2 Conditional radial modelling
Viewing the angular variable $\mathbf{W}$ as a ‘covariate’ for the radial variable $R$, inference for the SPAR model is analogous to a non-stationary POT analysis, for which many parametric and semi-parametric approaches have been proposed [chavez2005, Randell2016, Youngman2019, Zanini2020, Barlow2023]. Non-stationary POT can be performed using GP regression (modelling GP parameters as functions of covariates), which proceeds by first estimating the threshold function, $u(\mathbf{w})$, and then estimating the GP scale and shape parameter functions, $\sigma(\mathbf{w})$ and $\xi(\mathbf{w})$ respectively, via likelihood-based inference procedures. The choice of the functional forms for $u(\mathbf{w})$, $\sigma(\mathbf{w})$, and $\xi(\mathbf{w})$ determine the flexibility of the overall model. As the dimension $d$ grows, these mappings become increasingly complex, and so models that represent the functions via semi-parametric models (such as the splines used in initial work with the SPAR model [murphy2024inference] and other similar angular-radial approaches [simpson2024estimating, majumder2023semiparametric]) are unlikely to offer sufficient flexibility. Moreover, they become increasingly computationally-demanding to estimate for large $d$. Consequently, we adopt a deep learning approach, whereby the threshold and GP parameter functions are represented using artificial neural networks (ANN); for details on GP regression with deep learning, see [Richards2024]. Although the representation of the SPAR parameter functions via ANNs differs from previous approaches in this framework, the ‘loss function’ used for optimisation of the model, is the same. That is, we estimate the threshold and GP parameter functions that maximise the likelihood evaluated on a hold-out dataset, as described in Section 3.2.2. This work builds upon the approach of [murphy2024deep], who use deep learning to estimate the extremal dependence structure of random vectors on standardised marginal scales via a similar angular-radial decomposition. In contrast, our approach does not require marginal transformation, and is thus not subject to marginal estimation uncertainty.
@@ 3.2.1 Neural network representation of conditional radial parameters
Several recent approaches have used deep learning for GP regression; see, e.g. [wilson2022deepgpd, richards2023insights, cisneros2024deep, pasche2024neural]. In these approaches, ANNs are used to model the relationships between covariates and GP parameter and threshold functions. Our approach is analogous, with the covariates taken to be angles on the hypersphere. Note that the hypersphere is compact (i.e., closed and bounded), and that this is a desirable property for extrapolation in deep learning.
@@
Here, we design two neural network models; one for the threshold function $u(\mathbf{w})$ and one for the parameter vector $(\nu(\mathbf{w}),\xi(\mathbf{w}))$, where $\nu(\mathbf{w})=\sigma(\mathbf{w})(\xi(\mathbf{w})+1)$ is the modified scale parameter. Unlike $\sigma(\mathbf{w})$, the modified scale parameter $\nu(\mathbf{w})$ is orthogonal to $\xi(\mathbf{w})$, which helps to mitigate numerical instabilities during model fitting [cox1987parameter, chavez2005, tendijck2024practical]; [pasche2024neural] show that this is particularly helpful when estimating deep GP regression models. Both $u(\mathbf{w})$ and $(\nu(\mathbf{w}),\xi(\mathbf{w}))$ are modelled by multi-layer perceptrons (MLPs), which are a standard class of fully-connected ANN that compose multiple layers of ‘neurons’ [goodfellow2016deep]. Each neuron passes a linear combination of input variables through a nonlinear ‘activation function’, and the output is then passed to the subsequent layer; detailed discussions and illustrative figures can be found in [Richards2024], and an example schematic for an MLP representation of $(\nu(\mathbf{w}),\xi(\mathbf{w}))$ is presented in Figure 1. Inference for the parameter functions then involves estimating the linear coefficients (the ‘weights’ and ‘biases’) in each neuron of the corresponding MLP. Prior to inference, the architecture of the MLP must be defined, i.e., the number of hidden layers, denoted by $L$, the number of neurons in each hidden layer, denoted by $h_{1},\ldots,h_{L}$, and the type(s) of activation function(s). The resulting set of estimable parameters for the MLP contains all of the weights and biases in each hidden layer, as well as the final $(L+1)$-th layer; we denote this by $\mathcal{W}=\left\{\left(a^{l},b^{l}\right);l=1,\ldots,L+1\right\}$, with weights $a^{l}\in\mathbb{R}^{h_{l}\times h_{l-1}}$ and biases $b^{l}\in\mathbb{R}^{h_{l}}$. Note that the estimable sets of parameters differ between the MLPs for $u(\mathbf{w})$ and $(\nu(\mathbf{w}),\xi(\mathbf{w}))$; we denote their respective parameter sets by $\mathcal{W}_{u}$ and $\mathcal{W}_{(\nu,\xi)}$. For both MLPs, we take all hidden layer activation functions to be the rectified linear unit function, $\text{ReLU}(\mathbf{x})=(\max\{x_{1},0\},\max\{x_{2},0\},\dots)$. The final layers of the MLPs make use of an exponential transformation to ensure that the scale $\nu(\mathbf{w})$ and threshold $u(\mathbf{w})$ are strictly positive, that is, $\nu(\mathbf{w})>0,u(\mathbf{w})>0$ for all $\mathbf{w}$. For numerical stability, we also ensure that $\xi(\mathbf{w})$ satisfies $\xi(\mathbf{w})\in(-0.5,0.1)$ for all angles $\mathbf{w}$. Selection of the remaining tuning parameters is discussed in Section 3.2.3.
@@
Figure 1: Example schematic of an MLP model with $L=2$ hidden layers. The inputs are the components of the angle $\mathbf{w}=(w_{1},w_{2},...,w_{d})$ and the outputs are the GP parameter functions $(\nu(\mathbf{w}),\xi(\mathbf{w}))$.
@@ 3.2.2 Estimating the neural network parameters
To obtain estimates of the MLP parameter sets, $\mathcal{W}_{u}$ and $\mathcal{W}_{(\nu,\xi)}$, we optimise specified loss functions. Suppose that we have a set of radial and angular observations $\{(r_{i},\mathbf{w}_{i});i=1,...,n\}$. Recall from Equation (6) that the threshold $u(\mathbf{w})$ is taken to be the $1-\zeta$ quantile of $R|\mathbf{W}=\mathbf{w}$. We thus can estimate $u(\mathbf{w})$ (and its corresponding parameter set $\mathcal{W}_{u}$) using techniques from quantile regression [koenker2017handbook]. In this case, the most appropriate loss function for $u(\mathbf{w})$ is the tilted loss,
@@
After estimation of $u(\mathbf{w})$, we define $I_{u}:=\{i:r_{i}>u(\mathbf{w}_{i})\}$ as the set of indices of radial threshold exceedances. The MLP that defines the GP parameter functions can be considered as a ‘conditional density estimation network’, with the negative log-likelihood function used for optimisation; see, e.g. [neuneier1994estimation, rothfuss2019conditional]. In this case, we perform maximum likelihood estimation, and thus the loss function is given by
$\mathcal{L}_{GP}(\mathcal{W}_{(\nu,\xi)}):=-\sum_{i\in I_{u}}\log\left[f_{GP}\left(r_{i}-u(\mathbf{w}_{i});\xi(\mathbf{w}_{i}),\frac{\nu(\mathbf{w}_{i})}{\xi(\mathbf{w}_{i})+1}\right)\right].$ (10)
@@
As noted in Section 2, when the GP shape parameter $\xi(\mathbf{w})$ is negative, the distribution of $R|\mathbf{w}=\mathbf{w}$ has a finite upper endpoint. Training of a deep GP regression model which permits negative shape parameter values can be computationally troublesome; see discussion by [richards2022regression]. At a given angle $\mathbf{w}_{i}$, if $\xi(\mathbf{w}_{i})<0$ and the radial observation $r_{i}$ exceeds the upper endpoint, i.e., $r_{i}>u(\mathbf{w}_{i})-\sigma(\mathbf{w}_{i})/\xi(\mathbf{w}_{i})$, then the loss function in (10) will evaluate to a non-finite value. Consequently, the loss surface over which we optimise $\mathcal{W}_{(\nu,\xi)}$ is highly irregular, and iterative gradient descent methods (like ADAM) may have trouble finding global maxima, or may predict out-of-sample parameter estimates that are infeasible, i.e., the loss is non-finite. To circumvent these issues during training, we initialise the MLP to ensure that the shape parameter function $\xi(\mathbf{w})$ is non-negative for all angles $\mathbf{w}$; in this way, at the outset of the training procedure, the loss function is guaranteed to be finite for all $w_{i},i=1,\dots,n$. Then, if the gradient descent optimisation algorithm produces non-finite loss values during training, we restart training (from the last iteration with finite loss values) with a smaller learning rate. Note that the fully-trained MLP may still provide negative values of $\xi(\mathbf{w})$. This training procedure was found to produce reliable estimates in our application.

### v2

@@ 3.1 Angular modelling
Estimation of densities on the hypersphere $\mathbb{S}^{d-1}$ is part of a discipline known as directional statistics [mardia2000, ley2017modern]. The key difference from estimation of densities on $\mathbb{R}^{d}$ is that the surface of the hypersphere is periodic and bounded, and so distributions defined on $\mathbb{S}^{d-1}$ must conserve these constraints. Various parametric and non-parametric approaches have been developed for estimating densities on the hypersphere, which are directly analogous to approaches used in Euclidean space. These include kernel density (KD) estimation [hall1987kernel, bai1989kernel], mixture models [mooney2003fitting, fu2008modified, hornik2014movmf], and spline-based methods [ferreira2008directional]. See [Pewsey2021] for a recent review of non-parametric approaches, and [wessel2025comparison] for a comparison of various generative deep learning methods for modelling angular densities.
In keeping with the previous uses of SPAR in two dimensions, in this work we use a KD method for the angular modelling. The advantages of KD estimation are that it is simple to implement and fast to sample from. The main downside is that explicit calculation of the density can be slow for large sample sizes. However, for our purpose, simulation from the model is more important than explicit calculation of the density, so this drawback is less significant.
3.1.1 Kernel density estimation for spherical data
Suppose that $K(\mathbf{w};\bm{\mu},\kappa)$ is a density function on $\mathbb{S}^{d-1}$, with mean direction $\bm{\mu}\in\mathbb{S}^{d-1}$, bandwidth parameter $\kappa\geq 0$ that controls the concentration of the density around $\bm{\mu}$. Then the KD estimate of the angular density $f_{\mathbf{W}}$ from a sample of $n$ observations $\{\mathbf{w}_{i}\}_{i=1,...,n}\subset\mathbb{S}^{d-1}$ is given by
$\hat{f}_{\mathbf{W}}(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}K(\mathbf{w};\mathbf{w}_{i},\kappa),$ (10)
That is, the estimate is effectively a mixture of $n$ kernels, centred at each data point. From a modelling perspective, there are two key choices to make: the type of kernel to use, and the determination of the bandwidth.
3.1.2 Choice of kernel
The choice of kernel function has a relatively minor impact on the accuracy of the density estimate compared to the choice of bandwidth. While different kernels can slightly alter the shape of the density estimate, the bandwidth parameter, which controls the amount of smoothing, is the most crucial factor affecting the accuracy [scott2015multivariate]. Here we are mainly guided by computational considerations. The most common choice of kernel for KD models on the sphere is the von Mises-Fisher (vMF) distribution [fisher1953dispersion]. However, simulation from this distribution requires rejection sampling, which can be slow. Therefore, we opt to use the power spherical (PS) distribution [deCao2020psd]. The PS and vMF distributions are both rotationally symmetric distribution about the mean direction, and have similar shapes. However, simulation from the PS distribution is much faster, as discussed further in Appendix B. For $\mathbf{w}\in\mathbb{S}^{d-1}$ the density of the PS distribution is given by
$K(\mathbf{w};\bm{\mu},\kappa)=(4\pi)^{-\eta}\frac{\Gamma(2\eta+\kappa)}{\Gamma(\eta+\kappa)}\left[\tfrac{1}{2}(1+\mathbf{w}^{\top}\bm{\mu})\right]^{\kappa},$ (11)
where $\bm{\mu}\in\mathbb{S}^{d-1}$ is the mean direction, $\kappa\geq 0$ is the bandwidth parameter, $\eta=(d-1)/2$, and $\Gamma(\cdot)$ is the gamma function. The dot product of two unit vectors is the cosine of the angle between them. Therefore, the term $\mathbf{w}^{\top}\bm{\mu}\in[-1,1]$, defines the cosine of the arc length between $\mathbf{w}$ and $\bm{\mu}$, and plays an analogous role to the Euclidean distance $\|\mathbf{x}-\bm{\mu}\|_{2}\in[0,\infty)$ for isotropic kernels on $\mathbb{R}^{d}$.
For KD models, the bandwidth parameter $\kappa$ determines the balance between over-smoothing and over-fitting; it’s optimisation is discussed below. The PS distribution with $\kappa=0$ is the uniform distribution on $\mathbb{S}^{d-1}$. As $\kappa$ increases, the distribution becomes more concentrated about the mean direction. For larger sample sizes, using higher bandwidth parameters tends to result in a better resolution of complex distributional features. However, this can lead to numerical overflow problems when evaluating the gamma functions in (11), since $\Gamma(z)\sim\sqrt{2\pi z}(z/e)^{z}$ as $z\to\infty$. A more numerically stable expression is
$K(\mathbf{w};\mu,\kappa)=(4\pi)^{-\eta}\exp\left[\kappa\log(z)+\log\left(\Gamma(2\eta+\kappa)\right)-\log\left(\Gamma(\eta+\kappa)\right)\right],$
where $z=\tfrac{1}{2}(1+\mathbf{w}^{\top}\bm{\mu})\in[0,1]$ and hence $\log(z)\in(-\infty,0]$. We also have $\log(\Gamma(2\eta+\kappa))-\log(\Gamma(\eta+\kappa))\sim\eta\log(\kappa)$ as $\kappa\to\infty$. The log gamma function $\log(\Gamma(\cdot))$ can be computed directly in most software without having to evaluate the gamma function, thus avoiding overflow issues.
3.1.3 Bandwidth optimisation
There are various ways in which the bandwidth parameter for the KD model can be optimised, as discussed in [ley2017modern]. Here we use a cross-validation scheme to minimise the negative log-likelihood (NLL) for a hold-out sample. This is the most computationally expensive part of the method – for each test point, the kernel must be evaluated for each observation in the training sample. For a leave-one-out cross-validation scheme, this entails $(n-1)^{2}$ evaluations of the kernel:
$\mbox{NLL}(\kappa)=-\sum_{i=1}^{n}\log\left(\hat{f}_{\neq i}(\mathbf{w}_{i})\right),$ (12)
where $\hat{f}_{\neq i}(\mathbf{w}_{i})$ is the KD estimate (10) of the density from all observations except $\mathbf{w}_{i}$, evaluated at $\mathbf{w}_{i}$ (i.e., the predictive likelihood). However, as the NLL is effectively an expectation over $\mathbf{W}$, we can take a stochastic approach and estimate the NLL from a smaller sample of size $m$, chosen at random from the observations, so that we only need $(n-1)m$ kernel evaluations. This is similar to the approach used in stochastic gradient descent (see below) in machine learning, where a small batch of observations is used to approximate the true gradient of the cost function.
When observations are serially correlated, the estimate of $\hat{f}_{\neq i}(\mathbf{w}_{i})$ will be derived, in part, from observations that are correlated with $\mathbf{w}_{i}$, namely $\{...,\mathbf{w}_{i-2},\mathbf{w}_{i-1},\mathbf{w}_{i+1},\mathbf{w}_{i+2},...\}$. To obtain an unbiased estimate of the predictive likelihood using a cross-validation scheme, for each point $\mathbf{w}_{i}$ at which we want to calculate the predictive likelihood, we need to leave out observations around $\mathbf{w}_{i}$ which are correlated with it, say $\mathbf{w}_{i-k},...,\mathbf{w}_{i+k}$ for some fixed $k$.
[…]
Figure 1: Optimisation of bandwidth parameter $\kappa$ for kernel density estimate of the angular distribution, in terms of the minimum predictive negative log-likelihood.
3.1.4 Simulation from a kernel density estimate
Once the value of the bandwidth parameter has been optimised, it is straightforward to simulate from the KD model. As mentioned above, the KD model is effectively a mixture model, with one component corresponding to each observation, centred at the location of the observation. Therefore, to generate a random sample from the KD model requires two steps. First, we sample an observed angle $\mathbf{w}_{i}$ at random (i.e., a random integer $i\in\{1,...,n\}$). This corresponds to choosing one component from the mixture model at random. Secondly, we generate a random value from the kernel with mean direction $\mathbf{w}_{i}$ and bandwidth $\kappa$. As described in Appendix B, simulation from the PS kernel on $\mathbb{S}^{d-1}$ requires generating a beta random variable and $d-1$ standard normal random variables, and applying a rotation. Computationally this is very efficient. For the example presented in Section 4, generating a sample of size $2.7\times 10^{7}$ (100 times the observed sample size) took approximately 10 s on a laptop with specifications mentioned above.
@@ 3.2 Conditional radial modelling
Viewing the angular variable $\mathbf{W}$ as a ‘covariate’ for the radial variable $R$, inference for the SPAR model is analogous to a non-stationary univariate POT analysis, for which many parametric and semi-parametric approaches have been proposed [chavez2005, Randell2016, Youngman2019, Zanini2020, Barlow2023]. Non-stationary POT modelling can be performed using generalised Pareto (GP) regression (i.e., modelling GP parameters as functions of covariates), which proceeds by first estimating the threshold function, $u(\mathbf{w})$, via a quantile regression procedure, and then estimating the GP scale and shape parameter functions, $\sigma(\mathbf{w})$ and $\xi(\mathbf{w})$ respectively, via likelihood-based inference procedures. The choices of the functional forms for $u(\mathbf{w})$, $\sigma(\mathbf{w})$, and $\xi(\mathbf{w})$ determine the flexibility of the overall model. As the dimension $d$ grows, these mappings become increasingly complex, and models that represent the functions via semi-parametric models (such as the splines used in initial work with the SPAR model [murphy2024inference] and other similar angular-radial approaches [simpson2024estimating, majumder2023semiparametric]) become increasingly computationally-demanding to estimate. Consequently, we adopt a deep learning approach, whereby the threshold and GP parameter functions are represented using artificial neural networks (ANN). For details on GP regression with deep learning methods, see [Richards2024].
Although the representation of the GP parameter functions via ANNs differs from previous approaches in the SPAR framework, the ‘loss function’ used for optimisation of the model is the same. That is, we estimate the threshold and GP parameter functions that minimise the relevant loss functions evaluated on a hold-out dataset (in the case of the GP parameters, the loss function is the negative log-likelihood), as described in Section 3.2.2. This work builds upon the approach of [murphy2024deep], who use deep learning to estimate the extremal dependence structure of random vectors on standardised marginal scales via a similar angular-radial decomposition. In contrast, our approach does not require marginal transformation, and is thus not subject to marginal estimation uncertainty.
@@ 3.2.1 Neural network representation of conditional radial parameters
Several recent approaches have used deep learning for GP regression; see, e.g., [wilson2022deepgpd, richards2023insights, cisneros2024deep, pasche2024neural]. In these approaches, ANNs are used to model the relationships between covariates and GP parameter and threshold functions. Our approach is analogous, with the covariates taken to be angles on the hypersphere. Note that the hypersphere is compact (i.e., closed and bounded), and that this is a desirable property for extrapolation in deep learning.
@@
Here, we design two neural network models; one for the threshold function $u(\mathbf{w})$ and one for the parameter vector $(\nu(\mathbf{w}),\xi(\mathbf{w}))$, where $\nu(\mathbf{w})=\sigma(\mathbf{w})(\xi(\mathbf{w})+1)$ is the modified scale parameter. Unlike $\sigma(\mathbf{w})$, the modified scale parameter $\nu(\mathbf{w})$ is orthogonal to $\xi(\mathbf{w})$, which helps to mitigate numerical instabilities during model fitting [cox1987parameter, chavez2005, tendijck2024practical]; [pasche2024neural] show that this is particularly helpful when fitting deep GP regression models. Both $u(\mathbf{w})$ and $(\nu(\mathbf{w}),\xi(\mathbf{w}))$ are modelled by multi-layer perceptrons (MLPs), which are a standard class of fully-connected ANN that compose multiple layers of ‘neurons’ [goodfellow2016deep]. Each neuron passes a linear combination of input variables through a nonlinear ‘activation function’, and the output is then passed to the subsequent layer; detailed discussions and illustrative figures can be found in [Richards2024], and an example schematic for an MLP representation of $(\nu(\mathbf{w}),\xi(\mathbf{w}))$ is presented in Figure 2. Inference for the parameter functions involves estimating the linear coefficients (the ‘weights’ and ‘biases’) in each neuron of the corresponding MLP.
Prior to inference, the architecture of the MLP must be defined, i.e., the number of hidden layers, denoted by $L$, the number of neurons in each hidden layer, denoted by $h_{1},\ldots,h_{L}$, and the type(s) of activation function(s). The resulting set of estimable parameters for the MLP contains all of the weights and biases in each hidden layer, as well as the final $(L+1)$-th layer; we denote this by $\mathcal{W}:=\left\{\left(a^{l},b^{l}\right);l=1,\ldots,L+1\right\}$, with weights $a^{l}\in\mathbb{R}^{h_{l}\times h_{l-1}}$ and biases $b^{l}\in\mathbb{R}^{h_{l}}$. Note that the estimable sets of parameters differ between the MLPs for $u(\mathbf{w})$ and $(\nu(\mathbf{w}),\xi(\mathbf{w}))$; we denote their respective parameter sets by $\mathcal{W}_{u}$ and $\mathcal{W}_{(\nu,\xi)}$. For both MLPs, we take all hidden layer activation functions to be the rectified linear unit function, $\text{ReLU}(\mathbf{x})=(\max\{x_{1},0\},\max\{x_{2},0\},\dots)$. The final layers of the MLPs make use of an exponential transformation to ensure that the scale $\nu(\mathbf{w})$ and threshold $u(\mathbf{w})$ are strictly positive, i.e., $\nu(\mathbf{w})>0,u(\mathbf{w})>0$ for all $\mathbf{w}$. For numerical stability, we also ensure that $\xi(\mathbf{w})$ satisfies $\xi(\mathbf{w})\in(-0.5,0.1)$ for all angles $\mathbf{w}$. Selection of the remaining tuning parameters is discussed in Section 3.2.3.
@@
Figure 2: Example schematic of a multi-layered perceptron (MLP) model for the GP parameters, with $L=2$ hidden layers. The inputs are the components of the angle $\mathbf{w}=(w_{1},w_{2},...,w_{d})$ and the outputs are the GP parameter functions, $(\nu(\mathbf{w}),\xi(\mathbf{w}))$.
@@ 3.2.2 Estimating the neural network parameters
To obtain estimates of the MLP parameter sets, $\mathcal{W}_{u}$ and $\mathcal{W}_{(\nu,\xi)}$, we optimise specified loss functions. Suppose that we have a set of radial and angular observations $\{(r_{i},\mathbf{w}_{i});i=1,...,n\}$. Recall from Section 2.1 that the threshold $u(\mathbf{w})$ is taken to be the quantile of $R|(\mathbf{W}=\mathbf{w})$ at exceedance probability $\zeta$. We thus can estimate $u(\mathbf{w})$ (and its corresponding parameter set $\mathcal{W}_{u}$) using quantile regression techniques [koenker2017handbook]. In this case, an appropriate loss function for $u(\mathbf{w})$ is the tilted loss,
@@
After estimation of $u(\mathbf{w})$, we define $I_{u}:=\{i\in\{1,...,n\}:r_{i}>u(\mathbf{w}_{i})\}$ as the set of indices of radial threshold exceedances. The MLP that defines the GP parameter functions can be considered as a ‘conditional density estimation network’, with the negative log-likelihood function used for optimisation; see, e.g., [neuneier1994estimation, rothfuss2019conditional]. In this case, we perform maximum likelihood estimation, with loss
$\mathcal{L}_{GP}(\mathcal{W}_{(\nu,\xi)}):=-\sum_{i\in I_{u}}\log\left[f_{\rm GP}\left(r_{i}-u(\mathbf{w}_{i});\xi(\mathbf{w}_{i}),\frac{\nu(\mathbf{w}_{i})}{\xi(\mathbf{w}_{i})+1}\right)\right].$ (14)
@@
As noted in Section 2, when the GP shape parameter $\xi(\mathbf{w})$ is negative, the distribution of $R|(\mathbf{W}=\mathbf{w})$ has a finite upper endpoint. Training of a deep GP regression model which permits negative shape parameter values can be computationally troublesome; see discussion by [richards2022regression]. At a given angle $\mathbf{w}_{i}$, if $\xi(\mathbf{w}_{i})<0$ and the radial observation $r_{i}$ exceeds the upper endpoint, i.e., $r_{i}>u(\mathbf{w}_{i})-\sigma(\mathbf{w}_{i})/\xi(\mathbf{w}_{i})$, then the loss function in (14) will evaluate to a non-finite value. Consequently, the loss surface over which we optimise $\mathcal{W}_{(\nu,\xi)}$ is highly irregular, and iterative gradient descent methods (like ADAM) may have trouble finding global maxima, or may predict out-of-sample parameter estimates that are infeasible, i.e., the loss is non-finite. To circumvent these issues during training, we initialise the MLP to ensure that the shape parameter function $\xi(\mathbf{w})$ is non-negative for all angles $\mathbf{w}$; in this way, at the outset of the training procedure, the loss function is guaranteed to be finite for all $w_{i},i=1,\dots,n$. Then, if the gradient descent optimisation algorithm produces non-finite loss values during training, we restart training (from the last iteration with finite loss values) with a smaller learning rate. In our application we found that the fully-trained MLP provided negative values of $\xi(\mathbf{w})$ for all $\mathbf{w}\in\mathbb{S}^{d-1}$. We note that the issue of non-finite loss occurs in other likelihood-based inference procedures for GP regression, and is not a problem specific to MLPs.